# Rango: Adaptive Retrieval-Augmented Proving for Automated Software Verification

### v2

@@ III-A Proof Retriever
Note that information retrieval techniques other than BM-25 can be substituted into Rango’s proof retriever. In Section V-E, we explore the use of retrieval techniques that rely on neural networks, known as dense retrieval techniques.
@@ V Evaluation
    How does adapting proof retrieval to the current proof step contribute to Rango’s ability to synthesize proofs?
@@
    How does Rango’s rollout search compare to best-first search?
  • RQ7:
@@ V-C RQ2: Contribution of Proof and Lemma Retrievers
V-D RQ3: Retrieving at each Proof Step
Recall that Rango’s proof retrieval mechanism adapts to the current proof state by retrieving relevant proofs at every step in the proof search. To measure the effect of retrieving at each proof step, we compare Rango to a variant that only performs proof retrieval at the first step in the proof search. In other words, the variant retrieves proofs based on the theorem statement, and uses the same retrieved proofs throughout the search. We compare Rango with this variant on our ablation set and show the results in Table VI. We observe that Rango proves 35% more theorems by retrieving proofs at every proof step than it does by retrieving proofs solely at the beginning of the proof search. Thus, retrieving at every proof step is critical to Rango’s success.
TABLE VI: Retrieving at each Proof Step
Proof Retrieval Frequency Theorems Proven
Every proof step $\nicefrac{{150}}{{500}}=30.0\%$
Only at beginning of proof $\nicefrac{{111}}{{500}}=22.2\%$
Takeaway 3: By retrieving similar proofs at every proof step, rather than solely retrieving at the beginning of the proof, Rango proves 35% more theorems.
V-E RQ4: Comparing Retrieval Algorithms for Proof Retrieval
@@
TABLE VII: Proof Retrieval Variants
@@
Recall that a main requirement of the proof retriever is the ability to compare proof states. Rango uses the BM-25 algorithm to determine the similarity between proof states. Alternatively, to compare two proof states, Rango could use a neural network to compute a vector embedding that contains semantic information about each proof state. Then, the similarity between two proof states could be defined as the cosine similarity between the vector embeddings. We implemented this kind of proof retriever using the CodeBert 125M parameter model [23]. We compare the proof retriever used in Rango to a version that uses dense embeddings in Table VII. We also include another popular sparse retrieval algorithm, TF-IDF in Table VII. Because the difference in performance between BM-25 and TF-IDF is small, we ran this ablation on the entire CoqStoq benchmark instead of our ablation set to ensure the precision of our comparison. Rango proves $46\%$ more theorems than a variant that uses CodeBert embeddings to retrieve proofs. Thus, embeddings from CodeBert do not capture similarities between proof states that are relevant for proof retrieval. We also see that Rango proves $1\%$ more theorems than a variant that uses TF-IDF for proof retrieval.
Takeaway 4: Using the BM-25 sparse retrieval technique for proof retrieval led to 46% more proven theorems than using CodeBert dense embeddings and 1% more proven theorems than using another sparse retrieval technique, TF-IDF.
V-F RQ5: Rango with Naïve Retrieval Variant
TABLE VIII: Comparing Rango, Rango-PRE, and a Hybrid Approach
@@ V-F RQ5: Rango with Naïve Retrieval Variant
Rango uses a proof retriever and a lemma retriever to gather relevant context for a fine-tuned LLM to use when generating the next tactic in a proof. A naïve form of retrieving proofs and lemmas could use the lines directly preceding the theorem being proven as context to the LLM. We call this retrieval technique prefix retrieval. In this section, we explore Rango-PRE: a variant of Rango whose only retrieval mechanism is prefix retrieval. Table VIII shows Rango-PRE’s performance on the CoqStoq benchmark, and on our two post-cutoff projects. Note that while Rango proves more theorems than Rango-PRE, the margin is small. We know that Rango-PRE cannot retrieve context outside of the current file. However, we observed that it still performed well compared to Rango, which can retrieve proofs and lemmas throughout a project. This led us to the following hypothesis: when the required context is in close proximity to the current theorem, Rango-PRE is preferable since it presents this context to the LLM as it was originally written. However, when the required context is not in close proximity to the current theorem, Rango is preferable since it is able to retrieve proofs and lemmas across the project. Following this hypothesis, we created Rango-Hybrid to capitalize on advantages from both Rango and Rango-PRE. Rango-Hybrid simply alternates between rollouts using Rango and Rango-PRE. We can see that Rango-Hybrid leads to a $4\%$ increase in the number of theorems proven by Rango, and a $7\%$ increase in the number of theorems proven by Rango-PRE.
Takeaway 5: Rango proves more theorems than Rango-PRE. Rango and Rango-PRE can be combined into a hybrid search procedure that proves $4\%$ more theorems than Rango, and $7\%$ more theorems than Rango-PRE.
V-G RQ6: Searcher Variants
TABLE IX: Searcher Variants
@@ V-G RQ6: Searcher Variants
We compare Rango’s rollout search with these two best-first search configurations on our ablation set. We show the results of this comparison in Table IX. We can see from Table IX that Rango proves $6\%$ more theorems than the configuration using beam decoding and $20\%$ more theorems than the configuration using temperature sampling.
Takeaway 6: Rollout search is a simple, yet effective technique for synthesizing proofs. When used in Rango, it synthesizes $6\%$ more theorems than best-first search.
V-H RQ7: Understanding Rango’s Proofs
@@
V-H1 Kinds of Theorems Rango can Prove
@@
Takeaway 7.1: Like other proof synthesis tools, Rango’s success-rate has a strong relationship with the length of the human-written proof. We also found that the performance of Rango and other proof synthesis tools decreased for files with many dependencies.
V-H2 Qualities of Proofs Generated by Rango
TABLE X: Proof Length and Edit Distance between Machine-Generated and Human-Written Proofs
@@ V-H2 Qualities of Proofs Generated by Rango
To understand the qualities of proofs generated by Rango, we measure two attributes: proof length and edit distance to the human-written proof. We measure these attributes for Rango, Tactician and Proverbot. To calculate proof length, we first collected the $1,252$ theorems for which all three proof synthesis tools found proofs. Then, we calculated the number of tactics in each proof. We report the mean and median of these proof lengths in Table X. We also report the corresponding statistics for human-written proofs. Human-written proofs are shorter on average than proofs generated by all proof synthesis tools. Proofs generated by Rango and Tactician are of similar length, and are shorter on average than proofs generated by Proverbot.
We use string edit distance to the human-written proof as a measure of how “human-like” the proofs generated by Rango are. For each proof of the $1,252$ theorems where Rango, Tactician, and Proverbot all found proofs, we computed the string edit distance between the proofs found by the proof synthesis tools, and their corresponding human-written proofs. We report these edit distances in Table X. On average, Rango produces proofs that are 12 edit operations closer to human-written proofs than Tactician, and 21 edit operations closer to human-written proofs than Proverbot. A prior study on proof engineers found that they are constantly making repetitive repairs across their proofs due to changes to specifications or dependencies [71]. It is likely that automatically generated proofs that are in a similar style to the proof engineer’s hand-written proofs would be easier for them to repair and maintain, though future work should explore this.
Takeaway 7.2: On average, the proofs synthesized by Rango have similar or shorter lengths and smaller edit distances to human-written proofs than the proofs generated by other proof synthesis tools.
V-I Threats to Validity
@@ VI Related Work
Baldur [26] uses in-project information, using the lines preceding a theorem as context to its whole-proof-generation LLM. Section V-F showed that Rango can be instantiated with this retrieval mechanism, and that a hybrid search procedure between Rango and Rango-PRE leads to the best results on the CoqStoq benchmark. LEGO-Prover [84] builds a skill library of lemmas and theorems while proving so that it may retrieve new skills learned instead of relying on a fixed library.

### v3

@@ III-A Proof Retriever
Note that information retrieval techniques other than BM-25 can be substituted into Rango’s proof retriever. In Section V-D, we explore the use of retrieval techniques that rely on neural networks, known as dense retrieval techniques.
@@ V Evaluation
    How does Rango’s rollout search compare to best-first search?
@@ V-C RQ2: Contribution of Proof and Lemma Retrievers
V-D RQ3: Comparing Retrieval Algorithms for Proof Retrieval
@@
TABLE VI: Proof Retrieval Variants
@@
Recall that a main requirement of the proof retriever is the ability to compare proof states. Rango uses the BM-25 algorithm to determine the similarity between proof states. Alternatively, to compare two proof states, Rango could use a neural network to compute a vector embedding that contains semantic information about each proof state. Then, the similarity between two proof states could be defined as the cosine similarity between the vector embeddings. We implemented this kind of proof retriever using the CodeBert 125M parameter model [23]. We compare the proof retriever used in Rango to a version that uses dense embeddings in Table VI. We also include another popular sparse retrieval algorithm, TF-IDF in Table VI. Because the difference in performance between BM-25 and TF-IDF is small, we ran this ablation on the entire CoqStoq benchmark instead of our ablation set to ensure the precision of our comparison. Rango proves $46\%$ more theorems than a variant that uses CodeBert embeddings to retrieve proofs. Thus, embeddings from CodeBert do not capture similarities between proof states that are relevant for proof retrieval. We also see that Rango proves $1\%$ more theorems than a variant that uses TF-IDF for proof retrieval.
Takeaway 3: Using the BM-25 sparse retrieval technique for proof retrieval led to 46% more proven theorems than using CodeBert dense embeddings and 1% more proven theorems than using another sparse retrieval technique, TF-IDF.
V-E RQ4: Rango with Naïve Retrieval Variant
TABLE VII: Comparing Rango, Rango-PRE, and a Hybrid Approach
@@ V-E RQ4: Rango with Naïve Retrieval Variant
Rango uses a proof retriever and a lemma retriever to gather relevant context for a fine-tuned LLM to use when generating the next tactic in a proof. A naïve form of retrieving proofs and lemmas could use the lines directly preceding the theorem being proven as context to the LLM. We call this retrieval technique prefix retrieval. In this section, we explore Rango-PRE: a variant of Rango whose only retrieval mechanism is prefix retrieval. Table VII shows Rango-PRE’s performance on the CoqStoq benchmark, and on our two post-cutoff projects. Note that while Rango proves more theorems than Rango-PRE, the margin is small. We know that Rango-PRE cannot retrieve context outside of the current file. However, we observed that it still performed well compared to Rango, which can retrieve proofs and lemmas throughout a project. This led us to the following hypothesis: when the required context is in close proximity to the current theorem, Rango-PRE is preferable since it presents this context to the LLM as it was originally written. However, when the required context is not in close proximity to the current theorem, Rango is preferable since it is able to retrieve proofs and lemmas across the project. Following this hypothesis, we created Rango-Hybrid to capitalize on advantages from both Rango and Rango-PRE. Rango-Hybrid simply alternates between rollouts using Rango and Rango-PRE. We can see that Rango-Hybrid leads to a $4\%$ increase in the number of theorems proven by Rango, and a $7\%$ increase in the number of theorems proven by Rango-PRE.
Takeaway 4: Rango proves more theorems than Rango-PRE. Rango and Rango-PRE can be combined into a hybrid search procedure that proves $4\%$ more theorems than Rango, and $7\%$ more theorems than Rango-PRE.
V-F RQ5: Searcher Variants
TABLE VIII: Searcher Variants
@@ V-F RQ5: Searcher Variants
We compare Rango’s rollout search with these two best-first search configurations on our ablation set. We show the results of this comparison in Table VIII. We can see from Table VIII that Rango proves $6\%$ more theorems than the configuration using beam decoding and $20\%$ more theorems than the configuration using temperature sampling.
Takeaway 5: Rollout search is a simple, yet effective technique for synthesizing proofs. When used in Rango, it synthesizes $6\%$ more theorems than best-first search.
V-G RQ6: Understanding Rango’s Proofs
@@
V-G1 Kinds of Theorems Rango can Prove
@@
Takeaway 6.1: Like other proof synthesis tools, Rango’s success-rate has a strong relationship with the length of the human-written proof. We also found that the performance of Rango and other proof synthesis tools decreased for files with many dependencies.
V-G2 Qualities of Proofs Generated by Rango
TABLE IX: Proof Length and Edit Distance between Machine-Generated and Human-Written Proofs
@@ V-G2 Qualities of Proofs Generated by Rango
To understand the qualities of proofs generated by Rango, we measure two attributes: proof length and edit distance to the human-written proof. We measure these attributes for Rango, Tactician and Proverbot. To calculate proof length, we first collected the $1,252$ theorems for which all three proof synthesis tools found proofs. Then, we calculated the number of tactics in each proof. We report the mean and median of these proof lengths in Table IX. We also report the corresponding statistics for human-written proofs. Human-written proofs are shorter on average than proofs generated by all proof synthesis tools. Proofs generated by Rango and Tactician are of similar length, and are shorter on average than proofs generated by Proverbot.
We use string edit distance to the human-written proof as a measure of how “human-like” the proofs generated by Rango are. For each proof of the $1,252$ theorems where Rango, Tactician, and Proverbot all found proofs, we computed the string edit distance between the proofs found by the proof synthesis tools, and their corresponding human-written proofs. We report these edit distances in Table IX. On average, Rango produces proofs that are 12 edit operations closer to human-written proofs than Tactician, and 21 edit operations closer to human-written proofs than Proverbot. A prior study on proof engineers found that they are constantly making repetitive repairs across their proofs due to changes to specifications or dependencies [71]. It is likely that automatically generated proofs that are in a similar style to the proof engineer’s hand-written proofs would be easier for them to repair and maintain, though future work should explore this.
Takeaway 6.2: On average, the proofs synthesized by Rango have similar or shorter lengths and smaller edit distances to human-written proofs than the proofs generated by other proof synthesis tools.
V-H Threats to Validity
@@ VI Related Work
Baldur [26] uses in-project information, using the lines preceding a theorem as context to its whole-proof-generation LLM. Section V-E showed that Rango can be instantiated with this retrieval mechanism, and that a hybrid search procedure between Rango and Rango-PRE leads to the best results on the CoqStoq benchmark. LEGO-Prover [84] builds a skill library of lemmas and theorems while proving so that it may retrieve new skills learned instead of relying on a fixed library.